OpenStudy (anonymous):

Factor the following expression. 198q3r2 – 184q2r2 + 18qr2

jhonyy9 (jhonyy9):

33*2*3q^3r^2 -4*2*23q^2r^2 +3*3*2qr^2 = =2qr^2(3*11*3q^2 -4*23q +3*3) = =2qr^2(3*3(11q^2 +1)-4*23q)) = =2qr^2(9(11q^2 +1)-91q))